A ball is batted into the air and the ball found at a point 200 m distance horizontally in

5.0 seconds. Neglecting air resistance. Find the maximum height attain by the ball.


1
Expert's answer
2021-10-10T16:03:56-0400

The range "R=200m" and the time in air "T = 5s" are given as follows (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"R = \dfrac{u^2\sin2\theta}{g} \\
T = \dfrac{2u\sin\theta}{g}"


where "u" is the initial speed, "\theta" is the launch angle, "g = 9.81m\/s^2" is the gravitational acceleration.

The maximum height is given as follows:


"h = \dfrac{(u\sin\theta)^2}{2g}"

Expressing from the formula for time quantity "u\sin\theta" and substituting it into the expression for "h", obtain:


"u\sin\theta = \dfrac{gT}{2}\\
h = \dfrac{g^2T^2}{4\cdot 2g} = \dfrac{gT^2}{8}\\
h = \dfrac{9.81m/s^2\cdot (5s)^2}{8} \approx 31m"

Answer. 31 m.